# ICONS: Influence Consensus for Vision-Language Data Selection

### v3

@@ 2.4 Generalist: cross-task influence consensus
A similar aggregation approach is the one suggested in LESS xia2024less :
@@
An alternative approach which does not require directly comparing influence scores $\bar{I}_{k}$ across tasks $k$ is to leverage the relative rank of the training examples within each task. Concretely, we compute $\text{rank}_{k}({\bm{z}}_{i})$ for each example ${\bm{z}}_{i}$ relative to other examples for task $k$ according to their influence scores (higher influence scores correspond to lower rank). We can have a couple of options. First, we can select the training subset either using the Round Robin (RR) approach ivison2025large where we iterate over tasks and select the lowest-rank example which has not yet been selected to add to our set $\mathcal{S}_{\text{RR}}$. Alternatively, we can select the training subset $\mathcal{S}_{\text{MinRank}}$ such that all the examples within it have a low rank for some task $k$. Mathematically, albeit somewhat confusingly, this corresponds to:
@@
i.e., all examples that are not included in $\mathcal{S}$ would have high relative ranks $\text{rank}_{k}({\bm{z}}_{i})$ for all tasks $k$. However, this approach does not consider the potential interplay between tasks. Recall that in Eqn. 1 we aim to maximize sum of the relative performance across all tasks $k$; thus, if a training example is beneficial for multiple tasks, we may want to include it even at the expense of a lower-ranked example for a different task $k$. Thus, we introduce a simple consensus-based voting strategy that identifies training samples that consistently show a high influence score across various tasks. Concretely, we leverage the specialist training sets $\mathcal{S}_{k}$ as defined in §2.3 consisting of the $M$ highest-influence training examples for each task. We then select a combined training set as follows:
$S_{\text{ICONS}}=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=M}\sum_{k=1}^{K}\sum_{{\bm{z}}_{i}\in S}\mathbbm{1}[{\bm{z}}_{i}\in S_{k}]$ (6)
@@ 3 Experiments
In this section, we first discuss our experiment setup and evaluation benchmarks (§3.1). We then present our main results by comparing ICONS with the state-of-the-art methods (§3.2), followed by analysis of different selection strategies (§3.3). We further evaluate the transferability of ICONS3.4). Lastly, we provide analyze performance trends under different selection ratios (§3.5).
@@ 3.1 Evaluation test-bed
Datasets & model.  We apply ICONS on major visual instruction tuning (VIT) training datasets: LLaVA-665K liu2024improved , Cambrian-7M tong2024cambrian and Vision-Flan-186K visionflan . The majority of our analysis and ablation experiments are conducted on LLaVA-665K. For our experiments, we use the LLaVA-v1.5 model (liu2024improved, ) checkpoint after Stage 1 (pre-training for feature alignment) as defined in the original LLaVA training pipeline, with a default size of 7B parameters and LLaVA-665K unless otherwise specified. This checkpoint333 llava-v1.5-mlp2x-336px-pretrain-vicuna-7b-v1.5, which has no prior exposure to the visual instruction tuning data. corresponds to the model after training the projector but before any visual instruction tuning in Stage 2. Importantly, this model has not been exposed to the LLaVA-665K VIT dataset prior to the data selection process. In all experiments, we train the models for one epoch following the official finetuning hyperparameters using LoRA. More details on computation, including hardware specifications and runtime are in Appendix A.
@@
Target tasks.  We evaluate ICONS across diverse multimodal benchmarks (Appendix C, Tab. 5) that test different capabilities of vision-language models: 1) Multiple-choice understanding: MMBench (mmbench, ) and MME (mme, )  [^4] For MME, we focus on its perception section following coincide , which evaluates vision capablities. 2) Visual question answering: VQAv2 (vqav2, ), GQA (gqa, ), and VizWiz (vizwiz, ); 3) Text understanding in images: TextVQA (textvqa, ); 4) Scientific reasoning: ScienceQA (sqa, ); 5) Open-ended generation: LLaVA-W Bench (llava, ); 6) Factual consistency: POPE (pope, ).
@@
Baselines.  We compare our ICONS against several baselines, including random selection, CLIP-Score clip for measuring image-text alignment, EL2N el2n based on embedding L2 norms, and Perplexity perplexity using language model scores. We also compare against SemDeDup semdedup for semantic deduplication and D2-Pruning (d2, ) for distribution-aware pruning. Additional baselines include Self-Sup (selfsup, ) leveraging self-supervised signals, while Self-Filter (self-filter, ) and COINCIDE (coincide, ) are designed for vision-language data selection. We reference LLaVA-665K baseline results from COINCIDE coincide . Additionally, we compare with representation-based data selection baseline (RDS) ivison2025large ; xia2024less .
Method VQAv2 GQA VizWiz SQA-I TextVQA POPE MME MMBench LLaVA-W Rel. (%)
en cn Bench
Full 79.1 63.0 47.8 68.4 58.2 86.4 1476.9 66.1 58.9 67.9 100
Random 75.7 58.9 44.3 68.5 55.3 84.7 1483.0 62.2 54.8 65.0 95.8
CLIP-Score clip 73.4 51.4 43.0 65.0 54.7 85.3 1331.6 55.2 52.0 66.2 91.2
EL2N el2n 76.2 58.7 43.7 65.5 53.0 84.3 1439.5 53.2 47.4 64.9 92.0
Perplexity perplexity 75.8 57.0 47.8 65.1 52.8 82.6 1341.4 52.0 45.8 68.3 91.6
SemDeDup semdedup 74.2 54.5 46.9 65.8 55.5 84.7 1376.9 52.2 48.5 70.0 92.6
D2-Pruning d2 73.0 58.4 41.9 69.3 51.8 85.7 1391.2 65.7 57.6 63.9 94.8
Self-Sup selfsup 74.9 59.5 46.0 67.8 49.3 83.5 1335.9 61.4 53.8 63.3 93.4
Self-Filter self-filter 73.7 58.3 53.2 61.4 52.9 83.8 1306.2 48.8 45.3 64.9 90.9
COINCIDE coincide 76.5 59.8 46.8 69.2 55.6 86.1 1495.6 63.1 54.5 67.3 97.4
RDS ivison2025large ; xia2024less 75.1 57.9 48.6 68.0 54.9 86.3 1393.8 61.2 52.7 63.7 95.2
ICONS (ours) 76.3 60.7 50.1 70.8 55.6 87.5 1485.7 63.1 55.8 66.1 98.6
@@
Table 1: Selection results on LLaVA-665K. Performance comparison of different data selection approaches when trained on 20% of the LLaVA-665K dataset. The best and second best results for each benchmark are shown in bold and underlined, respectively. Our method ICONS achieves the highest overall Rel. (98.6%), consistently outperforming existing approaches including COINCIDE coincide (97.4%) and D2-Pruning d2 (94.8%).
@@ 3.2 Main results
ICONS (ours) 79.6 62.1 60.7 73.9 59.8 86.2 1,503.1 67.8 55.8 67.0 98.8
Table 2: Selection results on Vision-Flan-186K and Cambrian-7M. Performance comparison of different data selection approaches when trained on 20% of the Vision-Flan-186K visionflan and Cambrian-7M tong2024cambrian datasets. ICONS achieves strong performance (99.8% and 98.8% Rel. respectively) while using only 20% of the training data, significantly outperforming random selection which is one of the strongest baselines, and approaching full performance.
Cambrian-7M & Vision-Flan-186K selection.  We further provide results on Vision-Flan-186K and Cambrian-7M in Tab. 2. On Vision-Flan-186K, our method achieves near-full performance (99.8% Rel.) using just 37k samples, significantly outperforming random selection (91.6%) across tasks. Similarly, on Cambrian-7M, ICONS maintains a strong performance (98.8% Rel.) with 1,414k samples, while random selection achieves 95.4%. These results demonstrate that our approach scales effectively to both small and large datasets, consistently preserving model capabilities while drastically reducing the training data required.
Comparisons with representation-based data selection.  We compare our method against RDS (Representation-based Data Selection)ivison2025large ; xia2024less , a strong baseline in language-only instruction tuning. RDS computes training-validation similarity using final-layer representation of the last token in each sequence instead of gradients. For a fair comparison, we use the same influence matrix formulation (Eqn.2) and apply majority voting to reach influence consensus. Our method consistently outperforms RDS across all tasks, particularly those requiring perceptual grounding – e.g., higher scores on GQA (60.7 vs. 57.9), SQA-I (70.8 vs. 68.0), and MME (1485.7 vs. 1393.8). While RDS is effective in selecting large-scale text-only data (e.g. TULU-2/3 ivison2023camels ; lambert2024t ), its evaluation has largely focused on language-only tasks, where semantic similarity alone is often sufficient. In contrast, vision-language tasks demand alignment between modalities, where representation similarity is limited as it only reflects current embedding proximity and gradient-based approaches directly estimate each sample’s contribution to the validation loss. Our gradient-based approach directly estimates each sample’s impact on validation loss, capturing cross-modal training dynamics and prioritizing impactful training points. We further provide qualitative comparisons in Appendix §I.
@@ 3.3 Analysis of aggregation strategies
Ablations.  As introduced in §2.4, we explore multiple strategies for aggregating task-specific influence rankings into a single compact subset. We compare our majority voting approach (Vote) with different aggregation approaches for combining task-specific influence scores: 1) score-based methods (Merge, Max) and their normalized variants (Merge-GausNorm, Merge-SumNorm), 2) rank-based methods (Round Robin, MinRank). Our voting-based strategy outperforms all alternatives (Tab. 3), achieving the highest overall Rel. (98.6%). This shows that building a cross-task consensus via majority voting is a simple yet effective strategy for identifying consistently influential examples across tasks, without assuming calibration or comparability of scores.

### v4

@@ 3.4 Generalist: cross-task influence consensus
LESS [xia2024less] suggests a similar approach:
@@
An alternative approach which does not require directly comparing influence scores $\bar{I}_{k}$ across tasks $k$ is to leverage the relative rank of the training examples within each task. Concretely, we compute $\text{rank}_{k}({\bm{z}}_{i})$ for each example ${\bm{z}}_{i}$ relative to other examples for task $k$ according to their influence scores (higher influence scores correspond to lower rank). We can have a couple of options. First, we can select the training subset either using the Round Robin (RR) approach [ivison2025large] where we iterate over tasks and select the lowest-rank example which has not yet been selected to add to our set $\mathcal{S}_{\text{RR}}$. Alternatively, we can select the training subset $\mathcal{S}_{\text{MinRank}}$ such that all the examples within it have a low rank for some task $k$. Mathematically, albeit somewhat confusingly, this corresponds to:
@@
i.e., all examples that are not included in $\mathcal{S}$ would have high relative ranks $\text{rank}_{k}({\bm{z}}_{i})$ for all tasks $k$. However, this approach ignores the potential interplay between tasks. Since Eqn. 1 maximizes the sum of relative performance across tasks $k$, if a training example is beneficial for multiple tasks, we may want to include it even at the expense of lower-ranked examples for individual tasks. Thus, we introduce a consensus-based voting strategy that identifies training samples consistently showing high influence across tasks. For each task $k$, we set a threshold $\tau_{k}$ at the $((1-p)\times 100)$-th percentile of the influence score distribution, where $p$ represents the selection ratio ($p=0.2$ in our main experiments). A sample ${\bm{z}}_{i}$ is considered important for task $k$ if $\bar{I}_{k}({\bm{z}}_{i})\geq\tau_{k}$. We define the voting-based influence score as:
$I_{\text{vote}}({\bm{z}}_{i})=\sum_{k=1}^{K}\mathbbm{1}[\bar{I}_{k}({\bm{z}}_{i})\geq\tau_{k}],$ (6)
where $\mathbbm{1}[\cdot]$ is the indicator function. The final score ranges from 0 (no importance for any task) to $K$ (consistent importance for all tasks). Equivalently, we can define the threshold-based specialist sets as: $\mathcal{S}_{k}=\{{\bm{z}}_{i}\in\mathcal{D}:\bar{I}_{k}({\bm{z}}_{i})\geq\tau_{k}\}$. We then select the combined training set as:
$\mathcal{S}_{\text{ICONS}}$ $=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=M}\sum_{k=1}^{K}\sum_{{\bm{z}}_{i}\in S}\mathbbm{1}[{\bm{z}}_{i}\in S_{k}]$ (7)
$=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=M}\sum_{{\bm{z}}_{i}\in S}I_{\text{vote}}({\bm{z}}_{i})$ (8)
@@ 4 Experiments
We first discuss our experiment setup (§4.1), then compare ICONS with state-of-the-art methods (§4.2) and analyze different selection strategies (§4.3). We evaluate transferability to unseen tasks (§4.4), performance across selection ratios (§4.5), vote distributions (§4.6), and characteristics of low-vote samples (§4.7).
@@ 4.1 Evaluation test-bed
Datasets & model.  We apply ICONS on major VIT training datasets: LLaVA-665K [liu2024improved], Cambrian-7M [tong2024cambrian] and Vision-Flan-186K [visionflan]. The majority of our analysis and ablation experiments are conducted on LLaVA-665K. For our experiments, we use the LLaVA-v1.5 model [liu2024improved] checkpoint after Stage 1 (pre-training for feature alignment) as defined in the original LLaVA training pipeline, with a default size of 7B parameters and LLaVA-665K unless otherwise specified. This checkpoint333llava-v1.5-mlp2x-336px-pretrain-vicuna-7b-v1.5, which has no prior exposure to the visual instruction tuning data. corresponds to the model after training the projector but before any visual instruction tuning in Stage 2. Importantly, this model has not been exposed to the LLaVA-665K VIT dataset prior to the data selection process. In all experiments, we train the models for one epoch following the official finetuning hyperparameters using LoRA. More details on computation, including hardware specifications and runtime, are in Appendix §A.
@@
Target tasks.  We evaluate ICONS across diverse multimodal benchmarks (details are in Appendix §B, Tab. 6) that test different capabilities of vision-language models: 1) Multiple-choice understanding: MMBench [mmbench] and MME [mme]  [^4] For MME, we focus on its perception section following [coincide], which evaluates vision capabilities. 2) Visual question answering: VQAv2 [vqav2], GQA [gqa], and VizWiz [vizwiz]; 3) Text understanding in images: TextVQA [textvqa]; 4) Scientific reasoning: ScienceQA [sqa]; 5) Open-ended generation: LLaVA-W Bench [llava]; 6) Factual consistency: POPE [pope].
@@
Baselines.  We compare our ICONS against several baselines, including random selection, CLIP-Score [clip] for measuring image-text alignment, EL2N [el2n] based on embedding L2 norms, and Perplexity [perplexity] using language model scores. We also compare against SemDeDup [semdedup] for semantic deduplication and D2-Pruning [d2] for distribution-aware pruning. Additional baselines include Self-Sup [selfsup] leveraging self-supervised signals, while Self-Filter [self-filter] and COINCIDE [coincide] are designed for vision-language data selection. We reference LLaVA-665K baseline results from COINCIDE [coincide]. Additionally, we compare with representation-based data selection approach (RDS) [ivison2025large, xia2024less].
@@
Table 1: Selection results on LLaVA-665K. Performance comparison of different data selection approaches when trained on 20% of the LLaVA-665K dataset. Tasks are ordered by random selection’s relative performance. The best and second best results for each benchmark are shown in bold and underlined, respectively. Our method ICONS achieves the highest overall Rel. (98.6%), consistently outperforming existing approaches including COINCIDE [coincide] (97.4%) and D2-Pruning [d2] (94.8%).
Method MME SQA-I POPE VQAv2 LLaVA-W TextVQA MMBench GQA VizWiz Rel. (%)
Bench en cn
Full 1476.9 68.4 86.4 79.1 67.9 58.2 66.1 58.9 63.0 47.8 100
Random 1483.0 68.5 84.7 75.7 65.0 55.3 62.2 54.8 58.9 44.3 95.8
CLIP-Score [clip] 1331.6 65.0 85.3 73.4 66.2 54.7 55.2 52.0 51.4 43.0 91.2
EL2N [el2n] 1439.5 65.5 84.3 76.2 64.9 53.0 53.2 47.4 58.7 43.7 92.0
Perplexity [perplexity] 1341.4 65.1 82.6 75.8 68.3 52.8 52.0 45.8 57.0 47.8 91.6
SemDeDup [semdedup] 1376.9 65.8 84.7 74.2 70.0 55.5 52.2 48.5 54.5 46.9 92.6
D2-Pruning [d2] 1391.2 69.3 85.7 73.0 63.9 51.8 65.7 57.6 58.4 41.9 94.8
Self-Sup [selfsup] 1335.9 67.8 83.5 74.9 63.3 49.3 61.4 53.8 59.5 46.0 93.4
Self-Filter [self-filter] 1306.2 61.4 83.8 73.7 64.9 52.9 48.8 45.3 58.3 53.2 90.9
COINCIDE [coincide] 1495.6 69.2 86.1 76.5 67.3 55.6 63.1 54.5 59.8 46.8 97.4
RDS [ivison2025large, xia2024less] 1393.8 68.0 86.3 75.1 63.7 54.9 61.2 52.7 57.9 48.6 95.2
[…]
@@ 4.2 Main results
Table 2: Selection results on Vision-Flan-186K and Cambrian-7M. Performance comparison of different data selection approaches when trained on 20% of the Vision-Flan-186K [visionflan] and Cambrian-7M [tong2024cambrian] datasets. ICONS achieves strong performance (99.8% and 98.8% Rel. respectively) while using only 20% of the training data, significantly outperforming random selection which is one of the strongest baselines, and approaching full performance.
@@
[…]
Cambrian-7M & Vision-Flan-186K selection.  We provide results on Vision-Flan-186K [visionflan] and Cambrian-7M [tong2024cambrian] in Tab. 2. On Vision-Flan-186K, our method achieves near-full performance (99.8% Rel.) using just 37k samples, significantly outperforming random selection (91.6%). Similarly, on Cambrian-7M, ICONS maintains strong performance (98.8% Rel.) with 1,414k samples, while random selection achieves 95.4%. These results demonstrate our approach works effectively on both small and large datasets, consistently preserving model capabilities while drastically reducing training data.
Comparisons with representation-based data selection.  We compare against RDS (Representation-based Data Selection) [ivison2025large, xia2024less], a strong baseline in language-only data selection. RDS computes training-validation similarity using final-layer representation of the last token in each sequence instead of gradients. For a fair comparison, we use the same influence matrix formulation (Eqn. 2) and apply majority voting. Our method consistently outperforms RDS across all tasks, particularly those requiring perceptual grounding – e.g., higher scores on GQA (60.7 vs. 57.9), SQA-I (70.8 vs. 68.0), and MME (1485.7 vs. 1393.8). While RDS is effective for text-only data (e.g. TULU-2/3 [ivison2023camels, lambert2024t]), vision-language tasks demand alignment between modalities, where representation similarity is limited as it only reflects current embedding proximity, whereas gradient-based approaches directly estimate each sample’s contribution to validation loss, better capturing cross-modal training dynamics. We further provide qualitative comparisons in Appendix §F.
@@ 4.3 Analysis of aggregation strategies
Ablations.  As introduced in §3.4, we explore multiple strategies for aggregating task-specific influence rankings into a single compact subset. We compare our majority voting approach (Vote) with score-based methods (Merge, Max, and normalized variants (Merge-SumNorm, Merge-GausNorm)) and rank-based methods (Round Robin, MinRank). Our approach outperforms all alternatives (Tab. 3; full results in Appendix Tab. 5), achieving 98.6% Rel. Our majority voting effectively identifies consistently influential examples without requiring score calibration across tasks.
Table 3: Comparison of aggregation approaches. Relative performance of different influence aggregation methods when selecting 20% of the LLAVA-665K dataset on the ten tasks from Tables 1 and 2. Our Vote achieves the best performance (98.6% Rel.), outperforming score-based, their normalized variants, and rank-based baselines. See Appendix Tab. 5 for per-task results.
Type Score-based Normalized Score-based
Method Merge Max Merge-GausNorm Merge-SumNorm
[…]
Type Rank-based Vote-based
Method Round Robin MinRank ICONS (ours)
[…]
Limitations of score-based aggregation.  Score-based methods like Merge (Eqn. 3) and Max (Eqn. 4) assume calibrated influence scores across tasks, which is rarely the case. We observe substantial variation in influence score distributions across tasks with standard deviations spanning from $8.15\times 10^{-3}$ (MME) to $1.26\times 10^{-3}$ (VQAv2), indicating that influence scores for MME are much more spread out, while those for VQAv2 are tightly concentrated. Similarly, mean influence scores vary in both magnitude and sign: MME has a relatively high positive mean ($1.68\times 10^{-3}$), while tasks like POPE ($2.83\times 10^{-4}$) and GQA ($8.89\times 10^{-5}$) have negative means. These patterns show certain tasks have wider influence scores distributions, making a sample helpful for one task but neutral or harmful for another. Aggregating raw scores biases selection toward tasks with higher variance or skewed means. To address this calibration issue, we experiment with normalization: Merge-SumNorm, which rescales each task’s influence scores by dividing them by a task-specific normalization factor (i.e., sum), and Merge-GausNorm, which normalizes the scores using task-wise mean and standard deviation before averaging:
$I_{\text{Merge-SumNorm}}({\bm{z}}_{i})=\sum_{k=1}^{K}\frac{I_{k}({\bm{z}}_{i})}{\sum_{j}I_{k}({\bm{z}}_{j})}$
$I_{\text{Merge-GausNorm}}({\bm{z}}_{i})=\sum_{k=1}^{K}\frac{I_{k}({\bm{z}}_{i})-\mu_{k}}{\sigma_{k}}$
Limitations of rank-based aggregation.
Rank-based methods sidestep the challenge of comparing raw influence scores by focusing on within-task ranking. Round Robin selects samples by cycling through each task and picking the highest-scoring remaining sample for that task, ensuring balanced coverage. MinRank (Eqn. 5) selects samples that have the best minimum rank across all tasks, prioritizing examples that perform exceptionally well in at least one task regardless of their performance in others. Although these methods ensure balanced coverage across tasks, they can overfit to outlier tasks. This is particularly evident with LLaVA-W Bench [llava], which is an outlier in its influence ranking: both Round Robin and MinRank achieve relatively high scores on it (e.g. MinRank: 68.4) but at the cost of lower performance on all other tasks (Tab. 3). Rank-based selection can trade off overall efficacy on the mainstream tasks. In contrast, our Vote approach focuses on cross-task consensus rather than forcing equal representation, yielding better balance and higher overall performance (98.6% Rel.), showing the importance of identifying broadly influential examples rather than optimizing per-task rankings.
Figure 3: Unseen task generalization across methods. Performance on eight unseen benchmarks using 20% selected subsets, sorted by the average performance across 5 random selection runs (left: lower, right: higher). Shaded region shows mean and standard deviation across random runs. Our ICONS consistently outperforms all baselines. Full detailed results are in Appendix Tab. 7.
4.4 ICONS generalizes to unseen tasks
Data selected via ICONS demonstrate strong generalization on unseen benchmarks not used during data selection. In Fig. 3 and Appendix Tab. 7 we test LLaVA-ICONS-133K on diverse tasks including MMVet [mmvet], NaturalBench [naturalbench], AI2D [ai2d], ChartQA [chartqa], DocVQA [docvqa], InfoVQA [infographicvqa], RealWorldQA [xai2024grok] and CMMMU [cmmmu]. Notably, LLaVA-ICONS-133K achieves substantial improvements on challenging benchmarks (MMVet: 95.5%, AI2D: 97.3%) and even exceeds full dataset performance on harder tasks (CMMMU: 114.0%, RealWorldQA: 104.4%, InfoVQA: 103.8%, NaturalBench: 103.2%), demonstrating superior generalization to entirely unseen task formats and domains despite these tasks not being included in the selection process. Importantly, LLaVA-ICONS-133K outperforms random selection across all benchmarks. This suggests that our selection approach successfully captures fundamental visual-language understanding capabilities that transfer well across different task formats and domains. We further provide the cross-architecture generalization results in Appendix §C.
To understand how ICONS scales, we evaluate it across different selection ratios, progressively scaling the subset size from 5% to 90% of LLaVA-665K. As shown in Fig. 4, our results reveal several key patterns: First, ICONS shows particularly strong performance in the low-selection regime (5-20%), where identifying the most influential samples is crucial. Second, as the selection ratio increases, the performance gap between different methods gradually narrows. This convergence pattern is expected, as larger sample sizes naturally capture more of the dataset’s diversity and information. Despite this convergence trend, ICONS consistently outperforms all baselines across all selection ratios. Remarkably, it even surpasses full dataset performance at the 60% ratio, achieving over 102% relative score. One hypothesis is that ICONS can also effectively filter out potentially harmful or noisy training samples that might negatively impact model training, thereby surpassing the full training performance.
Figure 4: Different selection ratios. ICONS consistently outperforms all baselines across different selection ratios and remarkably exceeds 102% at a 60% selection ratio.
4.6 How does voting change with selection ratio?
We analyze vote distribution across selection ratios to understand task consensus patterns (Tab. 4, Appendix Fig. 6). At 5% selection, 72.1% of samples receive zero votes, indicating highly task-specific selection. As the ratio increases to 50% and 95%, the distribution shifts rightward with median votes rising to 4 and 5-7 respectively, and zero-vote samples decreasing to 8.3% and 1.2%. This rightward shift demonstrates that task consensus emerges as selection criteria relax, revealing samples with broader multi-task utility, while persistent zero-vote samples indicate universally low-quality or redundant data. These patterns validate our voting-based aggregation approach, which prioritizes samples with higher multi-task agreement. We further visualize zero-vote and one-vote samples in §4.7.
Table 4: Vote distribution across selection ratios. Vote distributions shift from task-specific selection at low ratios to broader consensus at high ratios.
[…]
5% 0.50 $\pm$0.98 0 7 72.1% 3 votes
20% 2.00 $\pm$1.56 2 7 28.4% 5 votes
50% 5.00 $\pm$2.21 5 7 8.3% 6 votes
90% 9.00 $\pm$1.84 9 10 1.2% 9 votes
95% 9.50 $\pm$1.42 10 10 0.8% 10 votes
4.7 What do low-vote training samples look like?
To understand what samples are filtered out, we examine zero-vote and 1-vote samples at a 90% selection ratio for LLaVA-665K, where the majority of samples receive votes, making the remaining low-vote samples particularly indicative of low-quality data. Zero-vote samples do not rank in the top percentile for any task, while 1-vote samples are selected by only a single task-specific specialist. As shown in Appendix Fig. 10, these low-vote samples exhibit characteristics that explain their low selection frequency: (1) subjective questions lacking objective ground truth (e.g., “Is the woman pretty?”), (2) trivial questions providing minimal visual learning signal (e.g., “Is the door open?” with answer “Yes”), and (3) redundant questions about the same image content. By filtering out such potentially harmful or redundant examples, ICONS focuses training on more informative and generalizable samples, and thus exceeds full-dataset performance.
5 Conclusion
We introduce ICONS, an influence consensus-based approach for visual instruction tuning data selection. It leverages gradient-based influence estimation and aggregates task-level influence scores through majority voting to select training samples broadly beneficial across tasks. ICONS addresses limitations of prior approaches by avoiding assumptions about score comparability across tasks and reducing sensitivity to outlier task rankings that bias selection. ICONS builds compact, high-impact datasets achieving 98.6% of full performance using only 20% of LLaVA-665K while generalizing well to unseen tasks and architectures. Beyond data selection, it provides a principled way to reason about data influence in multitask mixtures. We release LLaVA-ICONS-133K, Cambrian-ICONS-1.4M, and Vision-Flan-ICONS-37K as compact 20% subsets maintaining strong performance, hoping to inspire further exploration into data-efficient vision-language models.
Limitations.  Our approach’s main limitation is computational expense: computing gradients for large datasets is costly (Appendix §A), potentially constraining applicability to extremely large-scale data. To support the research community under resource constraints, we release the LLaVA-ICONS-133K dataset for iteration and model development.
Broader impact.  While our work does not directly imply negative impacts, it may indirectly propagate biases present in the original datasets. Positively, it enables more efficient and sustainable model development by reducing data redundancy and computational cost while maintaining performance. Continued discussion on these aspects remains critical for data-efficient vision-language models.
Acknowledgments.  This material is based upon work supported by the National Science Foundation under Grant No. 2107048 and No.2112562. Any opinions, findings, and conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. This work is also supported by the Singapore National Research Foundation and the National AI Group in the Singapore Ministry of Digital Development and Information under the AI Visiting Professorship Programme (award number AIVP-2024-001). We thank many people for their helpful discussion and feedback, listed in alphabetical order by last name: Allison Chen, Hee Seung Hwang, Hamish Ivison, Carlos E. Jimenez, Polina Kirichenko, Jaewoo Lee, Zhiqiu Lin, Tiffany Ling, Shengbang Tong, Ethan Tseng, Esin Tureci, Justin Wang, Zirui Wang.
[…]
Appendix A Computational complexity
A.1 Complexity analysis
A.2 Resource requirements
Table 5: Detailed aggregation approaches comparison for Tab. 3. Our proposed aggregation approach (Vote) consistently achieves the best overall performance (98.6% Rel.), outperforming both score-based (Merge, Max), their normalized variants (Merge-GausNorm, Merge-SumNorm) and rank-based (Round Robin, MinRank) baselines when selecting 20% of the LLAVA-665K dataset.